Please help me!!!
Multiply & Simlify
(3 + Square 2) (4 + Square 3)
Please help walk me through the steps of this problem...

Answer by stanbon(75887)

Multiply & Simlify
(3 + Square 2) (4 + Square 3)
----------------------------------
Use FOIL:
F: 3*4 = 12
O: 3*sqrt(3) = 3sqrt(3)
I: 4*sqrt(3) = 4sqrt(3)
L: sqrt(2)*sqrt(3) = sqrt(6)
------------------------------
Combine the terms to get:
= 12 + 7sqrt(3) + sqrt(6)
==============================
Cheers,
Stan H.